# Generalized Random Forests using Fixed-Point Trees

### v3

@@ 3.4 Pseudo-outcomes
Approximations $\tilde{\theta}_{C_{j}}$ of the form (9) and (3.3) offer an additional benefit: they enable criteria of the form (8) and (18) to be efficiently optimized by using a single multivariate CART split. A CART split performed with respect to vector-valued $\rho_{i}$ over a parent node $P$ produces child nodes $C_{1}$ and $C_{2}$ that minimize the following least-squares criterion:
@@
This equivalence is shown in Appendix B.1. GRF performs its splits by adopting gradient-based pseudo-outcomes, defined as
@@
such that the gradient-based approximation $\tilde{\theta}_{C_{j}}^{\texttt{grad}}$ of $\hat{\theta}_{C_{j}}$ can be written as
$\tilde{\theta}^{\texttt{grad}}_{C_{j}}=\hat{\theta}_{P}+\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i}^{\texttt{grad}}.$
In the case of fixed-point approximations, we define fixed-point pseudo-outcomes:
@@
$\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}\coloneqq-\eta\xi^{\top}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i}),\qquad\eta>0,$ (22)
@@
such that the fixed-point approximations $\tilde{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ of $\hat{\theta}_{C_{j}}$ can be written as
@@
$\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}=\hat{\theta}_{P}+\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}.$
@@
Substitute the above expression for $\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}$ into (18) to express fixed-point $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$-criterion in terms of the pseudo-outcomes:
@@
$\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\bar{\rho}^{~{}{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}}_{C_{1}}-\bar{\rho}^{~{}{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}}_{C_{2}}\right\rVert^{2},$
@@
where an analogous expression holds for gradient-based quantities $\tilde{\theta}_{C_{j}}^{\texttt{grad}}$, $\rho_{i}^{\texttt{grad}}$, and $\widetilde{\Delta}^{\texttt{grad}}$. We demonstrate in Appendix B.2 that maximizing the fixed-point criterion $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ is equivalent to maximizing the CART criterion (20), and extend this property to any $\widetilde{\Delta}$-style criterion based on a linear transformation of the parent scores $\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$.
@@
CART splits on pseudo-outcomes are computationally efficient. Given a parent node $P$, pseudo-outcomes of the form $\rho_{i}=-B\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$ are independent of the specific split $\{C_{1},C_{2}\}$ of $P$, for any matrix $B$ that is fixed with respect to the observations in $P$. This allows much of the computation for maximizing $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ to be done at the parent level, and avoids repeated calculations of $\tilde{\theta}_{C_{1}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ and $\tilde{\theta}_{C_{2}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ across the collection of candidate splits. After setting $P$ and computing $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$, the the value of $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ for the first candidate split requires $\mathcal{O}(n_{P})$ time, and the value for all other candidate splits of $P$ can be computed in $\mathcal{O}(1)$ time. While gradient-based pseudo-outcomes share this property, the use of fixed-point pseudo-outcomes eliminates the computational overhead and instability associated with estimating $A_{P}$, as discussed in Section 3.2.
@@
Fixed-point pseudo-outcomes defined in (22) include a scalar $\eta>0$. However, since tree splits are invariant to rescaling of the response, we can simply set $\eta=1$. With this adjustment, the fixed-point tree-growing procedure, referred to as the fixed-point tree algorithm, consists of two recursively applied steps:
@@ 3.5 Estimates of $\hat{\theta}(x)$ for Stage II
The fixed-point tree algorithm generates a single tree-based partition of $\mathcal{X}$. Repeating this process over subsamples of the training data yields a forest of trees, each specifying local leaf functions $L_{b}(x)$. These leaf functions define the local weight functions $\alpha_{i}(x)$ via (4), completing Stage I of GRF. The fixed-point tree training algorithm is described in Algorithm 1, while Algorithm 2 provides the pseudocode for the forest-wide Stage I procedure.
@@ 4 Theoretical Analysis
In this section, we provide a theoretical foundation for the GRF-FPT algorithm. For Stage I, Proposition 4.1 establishes an asymptotic equivalence between the approximate fixed-point criterion $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ and the target heterogeneity criterion $\Delta(C_{1},C_{2})$, providing a formal justification for the fixed-point tree algorithm.
@@
Suppose that Assumptions A.1 hold. Let $C_{1},C_{2}$ be candidate child nodes obtained from a fixed parent node $P$, and let $r\coloneqq\sup_{\{i:X_{i}\in P\}}\left\lVert X_{i}-x_{P}\right\rVert$, where $x_{P}$ denotes the center of mass with respect to the samples $X_{i}\in P$. Then, the fixed-point approximate criterion $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ is asymptotically equivalent to $\Delta(C_{1},C_{2})$ according to
$\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})=\Delta(C_{1},C_{2})+o_{P}\left(\max\left\{r^{2},\textstyle\frac{1}{n_{C_{1}}},\textstyle\frac{1}{n_{C_{2}}}\right\}\right),$ (24)
where $r>0$ is sufficiently small.
The proof of Proposition 4.1 can be found in Appendix B.3. The asymptotic rate (24) linking the fixed-point criterion $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ with the target criterion $\Delta(C_{1},C_{2})$ is the same rate established by Athey et al. (2019) between $\widetilde{\Delta}^{\texttt{grad}}(C_{1},C_{2})$ and $\Delta(C_{1},C_{2})$. This result provide a formal justification for the weight functions $\alpha_{i}(x)$ derived from the trees constructed in Stage I based on $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$.
For Stage II, Theorem 4.2 demonstrates the consistency of the GRF-FPT estimates $\hat{\theta}(x)$:
@@
Suppose that Assumptions A.1 hold, and let $(\hat{\theta}(x),\hat{\nu}(x))$ be estimates that solve (5) based on weights induced by a forest of trees grown under the fixed-point tree algorithm satisfying Specifications 1-5. Then, $(\hat{\theta}(x),\hat{\nu}(x))$ converges in probability to $(\theta^{*}(x),\nu^{*}(x))$.
The proof of Theorem 4.2 follows directly from Theorem 3 of Athey et al. (2019), which, under Assumptions A.1, establishes consistency for estimates $(\hat{\theta}(x),\hat{\nu}(x))$ that solve (5) with weights from a forest that satisfies Specifications 1-5. The forest specifications in Theorem 4.2 also apply to a forest trees grown under the fixed-point tree algorithm. Specifications 1-3 collectively impose mild boundary conditions on the splitting procedure. Meanwhile, Specification 4 requires that trees are trained on subsamples drawn without replacement (Biau et al., 2008; Scornet et al., 2015; Wager et al., 2014; Wager & Athey, 2018), and Specification 5 requires that trees must be grown using an additional subsample splitting mechanism known as honesty (Athey & Imbens, 2016; Biau, 2012; Denil et al., 2014). Appendix C.1 provides a detailed explanation of the subsampling and honest sample splitting procedure.
Finally, Theorem 4.3 establishes the asymptotic normality of the GRF-FPT estimates $\hat{\theta}(x)$:

### v4

@@ 3.4 Pseudo-outcomes
Approximations $\tilde{\theta}_{C_{j}}$ of the form (10) and (3.3) offer an additional benefit: they enable the $\widetilde{\Delta}$-criteria of the form (9) and (19) to be efficiently optimized through a single multivariate CART split. A CART split performed with respect to vector-valued responses $\rho_{i}\in\mathbb{R}^{K}$ over a parent node $P$ produces a split $(C_{1},C_{2})$ that minimizes the following least-squares criterion:
@@
The equivalence between the split that minimizes the least-squares CART criterion (20) and the split that maximizes (21) is shown in Appendix B.1.1. GRF performs its splits by adopting gradient-based pseudo-outcomes, defined as
@@
such that the gradient-based approximation $\tilde{\theta}_{C_{j}}^{\texttt{grad}}$ in (10) is equivalently written:
$\tilde{\theta}^{\texttt{grad}}_{C_{j}}=\hat{\theta}_{P}+\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i}^{\texttt{grad}}=\hat{\theta}_{P}+\bar{\rho}_{C_{j}}^{~{}{\texttt{grad}}}.$
In the case of fixed-point approximation, we define fixed-point pseudo-outcomes:
@@
$\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}\coloneqq-\eta\xi^{\top}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i}),\quad\eta\neq 0,$ (23)
@@
such that the fixed-point approximation $\tilde{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ in (3.3) is equivalently written as
@@
$\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}=\hat{\theta}_{P}+\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}=\hat{\theta}_{P}+\bar{\rho}_{C_{j}}^{~{}{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}}.$ (24)
@@
Substitute the above form of $\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{j}}$ into the $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$-criterion (19) to equivalently express the criterion in terms of the FPT pseudo-outcomes:
@@
$\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\bar{\rho}^{~{}{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}}_{C_{1}}-\bar{\rho}^{~{}{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}}_{C_{2}}\right\rVert^{2},$ (25)
@@
where an analogous equivalence holds for $\widetilde{\Delta}^{\texttt{grad}}$ in terms of the gradient-based pseudo-outcomes. We demonstrate in Lemma B.1 (in Appendix B.1.2) that maximizing the fixed-point criterion $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ is equivalent to maximizing the CART criterion (21), and extend this property to any $\widetilde{\Delta}$-style criterion induced by pseudo-outcomes that can be expressed as a split-independent linear transformation of the parent scores $\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$.
@@
Note that our method does not rely on iterative fixed-point procedures at all. Instead, it uses only a single step of fixed-point approximation to simplify the pseudo-outcomes. These simplified pseudo-outcomes are then passed directly to a standard CART algorithm for splitting. The numerical convergence of our method therefore relies solely on CART’s established and well-known stability, not on fixed-point iteration. CART splits on pseudo-outcomes are computationally efficient. Given a parent node $P$, the value $\rho_{i}=-B\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$ does not depend on a candidate split $(C_{1},C_{2})$ for any matrix $B$ that is fixed with respect to the parent. This allows much of the computation required to maximize $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ to be done at the parent level, and in particular avoids re-calculating the approximations $\tilde{\theta}_{C_{1}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ and $\tilde{\theta}_{C_{2}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ across the sequence of candidate splits. Once $P$ is fixed and $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ are computed, the value of $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ for the first candidate split requires $\mathcal{O}(n_{P})$ time, and the value for all other candidate splits of $P$ are queried in $\mathcal{O}(1)$ time. While gradient-based pseudo-outcomes share this property, the use of fixed-point pseudo-outcomes eliminates the computational overhead and instability associated with estimating $A_{P}$, as discussed in Section 3.2.
We show in Lemma B.2 (Appendix B.1.3) that choosing different values of $\eta$ does not change the outcome of the fixed-point splitting mechanism. Specifically, the optimal split identified by CART on pseudo-outcomes $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ of the form (23) does not depend on $\eta$. This can be heuristically understood by studying how the criterion changes as a function of the candidate splits. To illustrate, we consider a VCM model of the form (12) for bivariate regressors $W_{i}$, univariate $X_{i}\in[0,1]$, and scalar outcomes $Y_{i}$. A detailed summary of the settings is found in Appendix D.1. The sequence of valid candidate child nodes obtained by a split over univariate $X_{i}$ can be parameterized through scalar $t$ as $C_{1}(t)\coloneqq\{X_{i}\,:\,X_{i}\leq t\}$ and $C_{2}(t)\coloneqq\{X_{i}\,:\,X_{i}>t\}$. Let $\Delta(t)\coloneqq\Delta(C_{1}(t),C_{2}(t))$ denote the parameterized target criterion (8), and consider the behavior of $\Delta(t)$, $\widetilde{\Delta}^{\texttt{grad}}(t)$, and two fixed-point criteria $\widetilde{\Delta}_{1}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(t)$ and $\widetilde{\Delta}_{2}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(t)$ of the form (25) based on pseudo-outcomes with scale factors $\eta=1$ and $\eta=1/\sqrt{2}$, respectively. Figure 2 illustrates the different splitting criteria values plotted against the sequence of candidate splits. The visualization clearly shows that the criteria curves for $\Delta(t)$, $\widetilde{\Delta}^{\texttt{grad}}(t)$, and $\widetilde{\Delta}_{1}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(t)$ with $\eta=1$ are all very close to one another. Critically, the fixed-point criterion with $\eta=1/\sqrt{2}$, i.e. $\widetilde{\Delta}_{2}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(t)$, although scaled differently, still identifies the same maximizing split as $\widetilde{\Delta}_{1}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(t)$. This is because CART chooses a split based on a rank ordering of the criterion over all candidate splits. The absolute scale of the CART criterion does not matter, and it is only criterion rankings over the candidates that determines the optimal split. Therefore, choosing a different scalar $\eta$ does not change the outcome of the splitting process.
Based on the scale-invariance of our splitting criterion, we now detail the recursive procedure for growing our fixed-point trees pseudo-outcomes with $\eta=1$.
@@
The entire fixed-point tree-growing procedure recursively applies the following two steps on a given parent node $P$:
@@
Figure 2: Criterion values across candidate splits $(C_{1}(t),C_{2}(t))$ over threshold $t\in[0,1]$. The location of the optimal split under each criterion is given by the corresponding vertical line.
@@ 3.5 Estimates of $\hat{\theta}(x)$ for Stage II
The fixed-point tree algorithm generates a single tree-based partition of $\mathcal{X}$. Repeating this process over subsamples of the training data yields a forest of trees, each specifying local leaf functions $L_{b}(x)$. These leaf functions define the local weight functions $\alpha_{i}(x)$ via (4), completing Stage I of GRF. The full fixed-point tree training algorithm is described in Algorithm 1, while Algorithm 2 provides the pseudocode for the forest-wide Stage I procedure.
@@ 4 Theoretical Analysis
In this section, we provide a theoretical foundation for the GRF-FPT estimator $\hat{\theta}(x)$. For Stage I, Proposition 4.1 establishes an asymptotic equivalence between the FPT criterion and a weighted oracle criterion $\Delta_{V}(C_{1},C_{2})$ in (27), while Lemma 4.2 demonstrates that the Specifications A.2 are met by a forest based on the $\Delta_{V}$-criterion whenever they are met by a forest based on the $\Delta$-criterion. Assumptions A.1 and Specifications A.2 are the sufficient conditions for the consistency and asymptotic normality of $\hat{\theta}(x)$ in (5), and thus are used to formally justify the FPT algorithm as a mechanism for specifying an estimator of $\theta^{*}(x)$.
@@
Suppose Assumptions A.1 hold, and assume moreover Neyman orthogonal moment conditions (defined in Appendix A.4). Denote by $r\coloneqq\sup_{\{i:X_{i}\in P\}}\left\lVert X_{i}-x_{P}\right\rVert$ the radius of the parent $P$, where $x_{P}$ denotes the center of mass over $X_{i}\in P$. Let $V_{\theta\theta}(x_{P})$ denote the $\theta$-block of $V(x_{P})$ in (37). Denote by $\left\lVert\cdot\right\rVert_{V}$ the weighted Euclidean norm $\left\lVert z\right\rVert_{V}\coloneqq\left\lVert V_{\theta\theta}(x_{P})z\right\rVert_{2}=\sqrt{z^{\top}V_{\theta\theta}^{\top}(x_{P})V_{\theta\theta}(x_{P})z}$. Define the weighted oracle criterion $\Delta_{V}(C_{1},C_{2})$:
$\Delta_{V}(C_{1},C_{2})\coloneqq\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}\right\rVert_{V}^{2}.$ (27)
Then, treating the split as fixed with $r^{-2}\ll n_{C_{1}},n_{C_{2}}$ and sufficiently small $r>0$,
$\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})=\Delta_{V}(C_{1},C_{2})+o_{P}\left(r^{2},~{}\frac{1}{{n_{C_{1}}}},~{}\frac{1}{{n_{C_{2}}}}\right).$
Lemma 4.2.
Let $\mathcal{T}(\Delta)$ denote a tree whose splitting mechanism seeks splits that maximize $\Delta(C_{1},C_{2})$ defined in (8), and let $\mathcal{T}(\Delta_{V})$ denote a tree whose splitting mechanism seeks splits that maximize $\Delta_{V}(C_{1},C_{2})$ defined in (27). Suppose Assumptions A.1 hold and assume moreover that $\mathcal{T}(\Delta)$ is a tree that satisfies Specifications A.2. Then, $\mathcal{T}(\Delta_{V})$ satisfies Specifications A.2.
For Stage II, Theorem 4.3 establishes the consistency of the GRF-FPT estimator $\hat{\theta}(x)$:
@@
Suppose that Assumptions A.1 hold, and let $(\hat{\theta}(x),\hat{\nu}(x))$ be estimates that solve (5) based on weights induced by a forest of trees grown under the fixed-point tree algorithm satisfying Specifications A.2. Then, $(\hat{\theta}(x),\hat{\nu}(x))$ converges in probability to $(\theta^{*}(x),\nu^{*}(x))$.
The proof of Theorem 4.3 follows directly from Theorem 3 of Athey et al. (2019), which, under Assumptions A.1, establishes consistency for estimates $(\hat{\theta}(x),\hat{\nu}(x))$ that solve (5) with weights from a forest that satisfies Specifications 1-5. Thanks to Lemma 4.2, these forest specifications must also apply to a forest grown under the FPT mechanism. Specifications 1-3 collectively impose mild boundary conditions on the splitting procedure. Meanwhile, Specification 4 requires that trees are trained on subsamples drawn without replacement (Biau et al., 2008; Scornet et al., 2015; Wager et al., 2014; Wager & Athey, 2018), and Specification 5 requires that trees must be grown using an additional subsample splitting mechanism known as honesty (Athey & Imbens, 2016; Biau, 2012; Denil et al., 2014). Appendix C.1 provides a detailed explanation of the subsampling and honest sample splitting procedure.
Finally, Theorem 4.4 establishes the asymptotic normality of the GRF-FPT estimator $\hat{\theta}(x)$: